# Building a Data Dashboard for Magic: The Gathering: Initial Design Considerations Preprint — arXiv

### v1

@@ Abstract
{CCSXML}
<ccs2012> <concept> <concept_id>10003120.10003145.10003147.10010923</concept_id> <concept_desc>Human-centered computing Information visualization</concept_desc> <concept_significance>500</concept_significance> </concept> <concept> <concept_id>10003120.10003145.10011770</concept_id> <concept_desc>Human-centered computing Visualization design and evaluation methods</concept_desc> <concept_significance>500</concept_significance> </concept> </ccs2012>
@@
[…]
@@ 1 Introduction
Recent research in modern board and card games focuses on leveraging data-driven methods, artificial intelligence (AI) models, and visualization techniques to support decision-making, strategy development, and player understanding. Among these games, Magic: The Gathering (MTG) stands as the progenitor of the wider trading card game (TCG) genre as it is known today and the world’s largest tabletop collectible card game [ohm2023magic, zhu2023intelligence]. Specifically for MTG, some studies focus on generalized card representations, where multi-modal embeddings combine textual, numeric, visual, and usage-based features to capture card properties across expansions [bertram2024learning]. Other researchers tackle how game complexity has evolved through the years [magruder2022conservative] or card synergies [xue2025cardsynergy] using data visualizations. Another major area of research examines how AI agents and prediction models help drafting and human decision-making [ward2021ai, bertram2021predicting, zhu2023intelligence, alvin2021toward]. Traditional drafting is a complex process requiring players to balance immediate card value against long-term deck coherence. Beyond mechanics and data, perceptual and aesthetic factors also play a significant role. For instance, Kallabis et al. [kallabis2025deceptive] showed that visual styling can bias how players perceive card strength, even when objective properties remain constant. Finally, other studies focus on the gamer culture of MTG [trammell2010magic, falcao2021conservatism], predicting personality traits and color preferences [cui2025identity], or the cultural aspects of designing cards [fornazari2016gender, zanescu2022designing].
@@
Together, these studies highlight an emerging research space at the intersection of AI, visualization, and trading card games. However, we identified two critical research gaps in MTG related research. First, there is a lack of research on MTG’s most popular format, Commander. This format has become especially popular due to its large deck size, singleton rules, and focus on synergy-driven gameplay. This richness makes deck construction and strategy development both rewarding and challenging, as players must navigate vast card pools, hidden synergies, and shifting metagames. Second, although community-driven resources and APIs (e.g., Scryfall (www.scryfall.com) and MTGJSON (www.mtgjson.com)) and deck-building websites (e.g., Archidekt (www.archidekt.com) or Moxfield (www.moxfield.com)) provide comprehensive card data or data visualizations to track deck metrics, there is currently no research on visual analytics for MTG gameplay data, leaving players without interactive tools to explore strategic patterns or deck performance.
@@
In this paper, we report on the initial stages of a design study to develop a visualization dashboard for Commander. Our approach draws from design studies [sedlmair2012design] in visualization research, focusing on supporting players in understanding synergies, situating their decks within broader strategic contexts, and assessing their performance against other players. Specifically, we contribute with:
@@
  • A user-task analysis to identify requirements for a data visualization dashboard tailored to the Commander format.
@@
  • An evaluation of users’ comprehension and preferences through a structured user test of data visualizations.
@@ 2 Related Work
Within human-computer interaction (HCI) research, user-centered design [norman1986user] has become a well-established framework for developing systems that prioritize usability and user acceptance [fdis20099241]. This approach emphasizes integrating the needs, skills, and preferences of stakeholders into the design process [kopec2017livinglab, wilkinson2014applying]. Engaging users not only accelerates prototyping cycles [signoretti2015trip, ferreira2014elderly], but also enhances the overall quality of outcomes, often leading to more impactful and widely adopted solutions [lindsay2012engaging]. A specific branch of user-centered design, participatory design (PD), has been employed to uncover relevant requirements in case studies [frishberg2011interactive, kahng2017cti]. By fostering co-creation, it enables users to actively contribute to the identification of suitable design directions[lee2017steps]. Evidence shows that PD increases customer satisfaction and loyalty in private-sector contexts [grissemann2012customer], while in the public sector it has become central to driving social innovation [voorberg2015systematic].
@@ 2.1 Design Studies in Information Visualization
Recent work in visualization and interactive systems demonstrates the value of user-centered, multi-view, and explainable approaches for complex data domains (e.g., [wang2022extending, ruan2024violet, kandel2024pd, sarigai2025dciwebmapper]). dciWebMapper [sarigai2025dciwebmapper] provides a scalable framework for exploring high-dimensional geospatial data through coordinated, interactive views, illustrating how linked visualizations support pattern discovery. In healthcare, PD-Insighter [kandel2024pd] shows that participatory design can enable clinicians and patients to monitor daily behaviors effectively through temporal and spatial visualizations. For AI and complex model interpretation, Violet [ruan2024violet] visualizes quantum neural network activations to make opaque decision processes understandable, while Wang et al. [wang2022extending] apply user-centric explainable AI design to graph neural network outputs in drug repurposing, emphasizing domain-specific interpretability. A small subset of works focuses on investigating practical knowledge, identifying user challenges and visualization opportunities rather than prototype implementation to support comprehension and navigation within other domains (e.g., [li2023knowledge, wolf2025identifying]). Together, these studies highlight design principles that are directly applicable to building dashboards for complex domains such as trading card games: integrating multi-view coordination, supporting exploratory analysis, explaining complex relationships, and grounding the interface in user needs and domain expertise. Our work builds upon these findings to identify user requirements and propose data visualizations that directly support MTG analysis.
@@ 2.2 Data Visualizations for Magic: The Gathering
State-of-the-art research offers limited evidence of applying data visualization to explore card synergies, gameplay data, or card collections. Nevertheless, several online resources provide visual and analytical tools for MTG, ranging from community-driven dashboards to professional-grade platforms. For example, Vincenzo Galante’s Looker dashboard (www.github.com/VincenzoGalante/magic-the-gathering) presents a comprehensive analysis of MTG data through common visualization types, including horizontal bar charts for yearly card releases and pie charts for color distribution (Figure 1). Similarly, another exploratory analysis (www.colinberan.com/2023/06/11/exploratory-data-analysis-of-magic-the-gathering-cards) employs pie charts for color balance and vertical bar charts for card rarity. Online deck-building platforms adopt comparable visual conventions. For instance, Archidekt uses pie charts to represent color counts and bar charts to illustrate mana-value distributions (Figure 2).
Figure 1: Screenshot of the Galante’s dashboard.
Figure 2: Screenshot of Archidekt’s data visualizations for collections.
We found two resources that are aligned our objectives. First, Pacheco’s Commander dashboard (www.github.com/gabi-pacheco/MTG-Commander-Analysis) analyzes 53 Commander matches played between four players from June 2023 to August 2024 (Figure 3). This dataset captures deck usage, player performance, win conditions, and match outcomes, offering insights into player behavior and strategic tendencies. The dashboard also integrates a variety of visual forms (such as arc diagrams, area charts, and scatterplots) that can be interactively filtered by decks and players. A major limitation is that this project is custom made and does not allow other players to view their data. It also lacks interactive features and the ability to gather insights that are not explicit in the visualizations. Second, Playgroup.gg is a platform for tracking MTG Commander games, designed to capture game details and provide deep statistical insights for both casual and competitive players through extensive data visualization features. Key visualization features include (i) game and deck stats dashboards with charts for metrics such as win rates, average turn lengths, commander usage, power level estimates, and leaderboard positions and (ii) individual player dashboards featuring most-played decks, historical performance, and social stats such as rivalries and achievements (Figure 4). Although this project showcases the highest data diversity, the charts are spread across the website without an easy way to integrate them or allow for filtering. Color palettes for categorical encodings also hinder users’ ability to identify individual marks. Another reported limitation hinges on how challenging it is to find custom and granular stats visualizations.
@@
Figure 3: Screenshots of the dashboard developed by Pacheco.
@@
Figure 4: Screenshots of the dashboards developed by Playgroup.gg.
Together, these resources provide valuable reference points for developing our own visualization dashboard for Commander, highlighting the most common tasks, data dimensions, and chart types employed in existing practice. However, their limitations also reveal key opportunities for improvement. First, most existing dashboards focus on descriptive statistics and isolated charts rather than on facilitating comparative or exploratory analysis. This restricts users’ ability to identify deeper gameplay patterns such as synergies between commanders, temporal trends in performance, or contextual factors influencing win conditions. Second, visual encodings often rely on color as the primary differentiator, leading to accessibility issues and difficulties in distinguishing overlapping marks. Third, limited interactivity and lack of customization prevent users from tailoring visualizations to their own analytical goals or filtering data according to personal playgroup contexts. Finally, the fragmentation of visual components across different pages impedes holistic understanding and sensemaking. In the next section, we present our dashboard design methodology focused on addressing these issues while catering for player needs.
@@ 3 Methodology
In our study, we followed Sedlmair et al.’s structured approach to establishing effective collaborations. In particular, we went through stages from learn to design with a focus on understanding players’ needs and pain points regarding current data collection and analysis tools as well as which data visualizations may be adequate in this context to provide untapped insights of Commander gameplay. Throughout the process, we balanced usability with analytical depth, ensuring that the resulting dashboard provided actionable, context-sensitive insights while remaining approachable for players with varying levels of engagement, technical skill, and visualization literacy. In turn, this study forms the foundation for the subsequent design of the visualization dashboard. Experimental materials, code, and results are available at an online repository (https://osf.io/rsz9g/overview?view_only=e50cc8aff82e47c580bb422633599487).
@@ 4 Precondition Phase
This section presents the stages of learn, winnow, and cast.
@@ 4.1 Learn
As we mentioned, we focus on Commander (https://magic.wizards.com/en/formats/commander) (formerly known as Elder Dragon Highlander or EDH), one of the most popular and distinctive formats in MTG. Unlike competitive one-on-one formats, Commander emphasizes creativity, social interaction, and long-term strategy, making it particularly appealing to casual and community-oriented players. Each deck contains exactly 100 cards, and with the exception of basic lands, no duplicates are allowed. Central to the format is the choice of a Commander, which is usually a legendary creature or planeswalker, that defines the deck’s color identity through its mana symbols and shapes its overall strategy. Games are most often played with four players, each starting at 40 life points, which is double the usual starting total. The Commander itself is always available to cast from the Command Zone and can be replayed with an increasing two-mana tax each time it is removed. In addition to regular win conditions, the format introduces a unique rule called Commander Damage, where a player loses the game if they receive 21 combat damage from a single Commander over the course of the game. Commander is strongly defined by its synergy-driven deck-building. The singleton rule and the enormous card pool of more than 20,000 unique cards foster highly diverse and often thematic decks. Social dynamics, such as table politics, threat perception, and temporary alliances, further distinguish the format from competitive play. This combination of diversity, strategic depth, and strong community engagement has made Commander a cornerstone of modern MTG.

### v2

# 1]organization=Instituto Universitário de Lisboa (ISCTE-IUL), Business Research Unit (BRU-IUL), addressline=Av. das Forças Armadas, postcode=1649-026, state=Lisboa, country=Portugal

@@ Abstract
keywords:
@@
highlights
Commander players prioritize outcome-driven metrics
Canonical visualizations yield significantly higher comprehension
Players prefer contextually relevant playgroup comparisons over global population-level data
Establishment of practical design guidelines for gameplay dashboards
@@ 1 Introduction
Recent research in modern board and card games focuses on leveraging data-driven methods, artificial intelligence (AI) models, and Information Visualization (InfoVis) techniques to support decision-making, strategy development, and player understanding. Among these games, Magic: The Gathering (MTG) stands as the progenitor of the wider trading card game (TCG) genre as it is known today and the world’s largest tabletop collectible card game [ohm2023magic, zhu2023intelligence]. Specifically for MTG, some studies focus on generalized card representations, where multi-modal embeddings combine textual, numeric, visual, and usage-based features to capture card properties across expansions [bertram2024learning]. Other researchers tackle how game complexity has evolved through the years [magruder2022conservative] or card synergies [xue2025cardsynergy] using data visualizations. Another major area of research examines how AI agents and prediction models help drafting and human decision-making [ward2021ai, bertram2021predicting, zhu2023intelligence, alvin2021toward]. Traditional drafting is a complex process requiring players to balance immediate card value against long-term deck coherence. Beyond mechanics and data, perceptual and aesthetic factors also play a significant role. For instance, Kallabis et al. [kallabis2025deceptive] showed that visual styling can bias how players perceive card strength, even when objective properties remain constant. Finally, other studies focus on the gamer culture of MTG [trammell2010magic, falcao2021conservatism], predicting personality traits and color preferences [cui2025identity], or the cultural aspects of designing cards [fornazari2016gender, zanescu2022designing].
@@
Together, these studies highlight an emerging research space at the intersection of AI, visualization, and trading card games. However, we identified two critical research gaps in MTG related research. First, there is a lack of research on MTG’s most popular format, Commander. This format has become especially popular due to its large deck size, singleton rules, and focus on synergy-driven gameplay. This richness makes deck construction and strategy development both rewarding and challenging, as players must navigate vast card pools, hidden synergies, and shifting metagames. Second, although community-driven resources and APIs (e.g., Scryfall [scryfallScryfallMagic] and MTGJSON [mtgjsonMTGJSON]) and deck-building websites (e.g., Archidekt [archidektDeckBuilder] or Moxfield [moxfieldMoxfield]) provide comprehensive card data or data visualizations to track deck metrics, there is currently no research on visual analytics for MTG gameplay data, leaving players without interactive tools to explore strategic patterns or deck performance. Therefore, we believe that a visualization dashboard enabling players to explore the game’s rich dataset will enhance the visualization community by demonstrating effective visual idioms and task designs, while also improving players’ awareness of gameplay mechanics in this complex game.
@@
In this paper, we report on the initial stages of a design study to develop a visualization dashboard for Commander. Our approach draws from design studies [sedlmair2012design] in visualization research, focusing on supporting players in understanding synergies, situating their decks within broader strategic contexts, and assessing their performance against other players. Specifically, we contribute to human-computer interaction (HCI) evaluation methods with:
@@
  • A user-task analysis to identify requirements for a data visualization dashboard tailored to the Commander format of MTG.
@@
  • An evaluation of users’ comprehension and preferences through a structured user test of data information visualizations.
@@ 2 Related Work
Within HCI research, user-centered design [norman1986user] has become a well-established framework for developing systems that prioritize usability and user acceptance [fdis20099241]. This approach emphasizes integrating the needs, skills, and preferences of stakeholders into the design process [kopec2017livinglab, wilkinson2014applying]. Engaging users not only accelerates prototyping cycles [signoretti2015trip, ferreira2014elderly], but also enhances the overall quality of outcomes, often leading to more impactful and widely adopted solutions [lindsay2012engaging]. A specific branch of user-centered design, participatory design (PD), has been employed to uncover relevant requirements in case studies [frishberg2011interactive, kahng2017cti]. By fostering co-creation, it enables users to actively contribute to the identification of suitable design directions[lee2017steps]. Evidence shows that PD increases customer satisfaction and loyalty in private-sector contexts [grissemann2012customer], while in the public sector it has become central to driving social innovation [voorberg2015systematic]. In this section, we begin by explaining which design studies frameworks are relevant within the HCI community, to motivate why we applied one of them in the development of our proposal. We then review the current state of the art in game visual analytics, finishing with a specification for MTG.
@@ 2.1 Design Studies in Information Visualization
Recent work in visualization and interactive systems demonstrates the value of user-centered, multi-view, and explainable approaches for complex data domains (e.g., [wang2022extending, ruan2024violet, kandel2024pd, sarigai2025dciwebmapper]). dciWebMapper [sarigai2025dciwebmapper] provides a scalable framework for exploring high-dimensional geospatial data through coordinated, interactive views, illustrating how linked visualizations support pattern discovery. In healthcare, PD-Insighter [kandel2024pd] shows that participatory design can enable clinicians and patients to monitor daily behaviors effectively through temporal and spatial visualizations. For AI and complex model interpretation, Violet [ruan2024violet] visualizes quantum neural network activations to make opaque decision processes understandable, while Wang et al. [wang2022extending] apply user-centric explainable AI design to graph neural network outputs in drug repurposing, emphasizing domain-specific interpretability. A small subset of works focuses on investigating practical knowledge, identifying user challenges and visualization opportunities rather than prototype implementation to support comprehension and navigation within other domains (e.g., [li2023knowledge, wolf2025identifying]). Together, these studies highlight design principles that are directly applicable to building dashboards for complex domains such as trading card games: integrating multi-view coordination, supporting exploratory analysis, explaining complex relationships, and grounding the interface in user needs and domain expertise. Our work builds upon these findings to identify user requirements and propose visualizations that directly support MTG analysis.
2.2 Data Visualizations for Games
Game analytics visualization has emerged as a critical field over the past decade, driven by the gaming industry’s evolution into a highly competitive market where understanding player behavior is essential for improving game quality and player experience [Wallner2014]. The prevalence of internet-enabled gaming devices now enables developers to remotely monitor every aspect of gameplay, accumulating vast amounts of player-game interaction data that has become viable for guiding decision-making throughout the design process [Wallner2015]. This growing interest in game telemetry has led to the emergence of game analytics as a distinct field concerned with discovering and communicating meaningful patterns in gaming data [Wallner2018]. The field heavily relies on visualization techniques to assist developers, researchers, and players in understanding and exploring complex, multimodal datasets [Bowman2012].
The most commonly employed visualization methods include fundamental charts, heatmaps, and scatterplots for statistical analysis of aggregated data, though these approaches often neglect sequential information across player behavior [Teng2024]. Heatmaps have proven particularly popular for visualizing single metrics like player deaths and spatial distribution, allowing players to reenact game courses and develop new strategies [Kriglstein2014, Moura2011]. More sophisticated approaches include node-edge graphs for revealing individual and aggregated player processes [Teng2023], hexbin maps with visual glyphs for multivariate information [Wallner2020], and animated maps for spatio-temporal analysis in games like MOBAs [Goncalves2018]. Specialized systems like PLATO integrate multiple techniques, including subgraph matching, pathfinding, and clustering for comprehensive gameplay analysis [Wallner2014], while novel approaches explore play-graphs for formal description of gameplay data [Wallner2013] and glyph-based visualizations for understanding problem-solving strategies [Nguyen2021].
Despite significant advances, the field faces ongoing challenges, including the dynamic nature and high dimensionality of game data, scalability issues with individual player trace visualization, and the need for customization across different stakeholder needs [Wallner2018, Teng2023]. User studies have provided valuable insights into visualization effectiveness, with research showing that while heatmaps excel at detecting hotspots, cluster representations prove superior for comparing variables and uncovering relationships [Kriglstein2014]. Educational applications have driven innovation in feedback visualization, exploring metaphorical and emotionally-engaging approaches beyond traditional charts and graphs [Canossa2022]. The field continues to evolve with developments in real-time analysis [Kozowski2020], iterative segmentation approaches for managing complex visualizations [Teng2023], and specialized tools for retrospective battle analysis in competitive gaming [Wallner2016].
MTG represents an ideal testbed for advancing game analytics visualization research, addressing critical gaps identified in current literature. While existing work focuses heavily on action-oriented games and aggregated statistical analysis, researchers emphasize the urgent need for visualization approaches that can capture “strategies or intent” and sequential decision-makin processes that current methods end up neglecting [Teng2024].
@@ 2.3 Data Visualizations for Magic: The Gathering
State-of-the-art research offers limited evidence of applying data visualization to explore card synergies, gameplay data, or card collections. In fact, looking at current literature on TCG, there are not studies focused on visualizing gameplay data. Nevertheless, several online resources provide visual and analytical tools for MTG, ranging from community-driven dashboards to professional-grade platforms. For example, Vincenzo Galante’s Looker dashboard [githubGitHubVincenzoGalantemagicthegathering] presents a comprehensive analysis of MTG data through common visualization types, including horizontal bar charts for yearly card releases and pie charts for color distribution (Figure 1). Similarly, another exploratory analysis [beran] employs pie charts for color balance and vertical bar charts for card rarity. Online deck-building platforms adopt comparable visual conventions. For instance, Archidekt uses pie charts to represent color counts and bar charts to illustrate mana-value distributions (Figure 2).
Figure 1: Screenshot of the Galante’s dashboard [githubGitHubVincenzoGalantemagicthegathering], a comprehensive analysis of MTG data through common visualization types.
Figure 2: Screenshot of Archidekt’s data visualizations for collections [archidektDeckBuilder] , where pie charts are used to represent color counts and bar charts to illustrate mana-value distributions.
We found two resources that are aligned our objectives. First, Pacheco’s Commander dashboard [githubGitHubGabipachecoMTGCommanderAnalysis] analyzes 53 Commander matches played between four players from June 2023 to August 2024 (Figure 3). This dataset captures deck usage, player performance, win conditions, and match outcomes, offering insights into player behavior and strategic tendencies. The dashboard also integrates a variety of visual forms (such as arc diagrams, area charts, and scatterplots) that can be interactively filtered by decks and players. A major limitation is that this project is custom made and does not allow other players to view their data. It also lacks interactive features and the ability to gather insights that are not explicit in the visualizations. Second, Playgroup.gg is a platform for tracking MTG Commander games, designed to capture game details and provide deep statistical insights for both casual and competitive players through extensive data visualization features. Key visualization features include (i) game and deck stats dashboards with charts for metrics such as win rates, average turn lengths, commander usage, power level estimates, and leaderboard positions and (ii) individual player dashboards featuring most-played decks, historical performance, and social stats such as rivalries and achievements (subsection 4.3). Although this project showcases the highest data diversity, the charts are spread across the website without an easy way to integrate them or allow for filtering. Color palettes for categorical encodings also hinder users’ ability to identify individual marks. Another reported limitation hinges on how challenging it is to find custom and granular stats visualizations.
@@
Figure 3: Screenshots of the dashboard developed by Pacheco [githubGitHubGabipachecoMTGCommanderAnalysis]. This dataset captures deck usage, player performance, win conditions, and match outcomes, providing insights into player behavior and strategic tendencies. The accompanying dashboard presents this data through interactive visualizations, including arc diagrams, area charts, and scatterplots, that can be filtered by deck or player.
@@
Figure 4: Screenshots of the dashboards developed by Playgroup.gg [playgroug]. Key visualization features include game and deck stats dashboards with charts for metrics such as win rates, average turn lengths, commander usage, power level estimates, and leaderboard positions. Individual player dashboards display most-played decks, historical performance, and social stats, including rivalries and achievements.
Together, these resources provide valuable reference points for developing our own visualization dashboard for Commander, highlighting the most common tasks, data dimensions, and chart types employed in existing practice. However, their limitations also reveal key opportunities for improvement. First, from a research perspective, most dashboards we identified lack support from scientific peer-review pipelines. They were developed by the community as solutions to specific problems. In contrast, we applied a formal design study framework. Second, most existing dashboards focus on descriptive statistics and isolated charts rather than on facilitating comparative or exploratory analysis. This restricts users’ ability to identify deeper gameplay patterns such as synergies between commanders, temporal trends in performance, or contextual factors influencing win conditions. Third, visual encodings often rely on color as the primary differentiator, leading to accessibility issues and difficulties in distinguishing overlapping marks. Forth, limited interactivity and lack of customization prevent users from tailoring visualizations to their own analytical goals or filtering data according to personal playgroup contexts. Finally, the fragmentation of visual components across different pages impedes holistic understanding and sensemaking. In the next section, we present our dashboard design methodology focused on addressing these issues while catering to player needs.
@@ 3 Methodology
In our study, we followed Sedlmair et al.’s structured approach to establishing effective collaborations. In particular, we went through stages from learn to design with a focus on understanding players’ needs and pain points regarding current data collection and analysis tools as well as which data visualizations may be adequate in this context to provide untapped insights of Commander gameplay. We conclude with the reflect phase, where we discuss our findings in relation to the design phase and outline how future work can build on our results to complete the pipeline through the implementation and deployment phases. Throughout the process, we balanced usability with analytical depth, ensuring that the resulting dashboard provided actionable, context-sensitive insights while remaining approachable for players with varying levels of engagement, technical skill, and visualization literacy. In turn, this study forms the foundation for the subsequent design of the visualization dashboard. Experimental materials, code, and results are available at an online repository  [^1] https://osf.io/rsz9g/overview?view_only=e50cc8aff82e47c580bb422633599487.
@@ 4 Precondition Phase
The first phase of the design study methodology by Sedlmair et al. [sedlmair2012design] is the precondition phase, which comprises three stages: learn, winnow, and cast.
@@ 4.1 Learn
As we mentioned, we focus on Commander [wizardsCommanderFormat] (formerly known as Elder Dragon Highlander or EDH), one of the most popular and distinctive formats in MTG. Unlike competitive one-on-one formats, Commander emphasizes creativity, social interaction, and long-term strategy, making it particularly appealing to casual and community-oriented players. Each deck contains exactly 100 cards, and with the exception of basic lands, no duplicates are allowed. Central to the format is the choice of a Commander, which is usually a legendary creature or planeswalker, that defines the deck’s color identity through its mana symbols and shapes its overall strategy. Games are most often played with four players, each starting at 40 life points, which is double the usual starting total. The Commander itself is always available to cast from the Command Zone and can be replayed with an increasing two-mana tax each time it is removed. In addition to regular win conditions, the format introduces a unique rule called Commander Damage, where a player loses the game if they receive 21 combat damage from a single Commander over the course of the game. Commander is strongly defined by its synergy-driven deck-building. The singleton rule and the enormous card pool of more than 20,000 unique cards foster highly diverse and often thematic decks. Social dynamics, such as table politics, threat perception, and temporary alliances, further distinguish the format from competitive play. This combination of diversity, strategic depth, and strong community engagement has made Commander a cornerstone of modern MTG.